Find the equation of the cylinder whose base is the circle x^2 + y^2 = 4, z = 0 and the axis is x= y/2 = z.
1
Expert's answer
2019-05-30T05:09:51-0400

All sections of the cylinder by planes with equation z = z0 are circles of radius 2 and center in the point of intersection of this plane and axis of the cylinder.

Equations of a circle with center (x0, y0) and radius r is


"(x-x_0)^2+(y-y_0)^2=r^2"

substituting x0 = z, y0 = 2z and r = 2 into this equation gives equation of the cylinder


"(x-z)^2+(y-2z)^2=4""x^2-2xz+z^2+y^2-4yz+4z^2=4"

"x^2+y^2+5z^2-2xz-4yz-4=0"

Answer: x2 + y2 +5z2 - 2xz - 4yz - 4 = 0.